OpenStudy (anonymous):

d/dx y=cos(1/x)/(1+x^2)

OpenStudy (nikvist):

\[y=\frac{\cos{(1/x)}}{1+x^2}\]\[y'=\frac{(\cos{(1/x))'(1+x^2)}-\cos{(1/x)(1+x^2)'}}{(1+x^2)^2}=\]\[=\frac{-\sin{(1/x)(-1/x^2)(1+x^2)}-2x\cos{(1/x)}}{(1+x^2)^2}=\]\[=\frac{(1+x^2)\sin{(1/x)}-2x^3\cos{(1/x)}}{(x+x^3)^2}\]